OpenStudy (anonymous):

If k(x) = 5x – 6, which expression is equivalent to (k + k)(4)?

OpenStudy (anonymous):

(k+k)x = 5x-6 +5x-6 (k+k)x = 10x-12 put x = 4 now

OpenStudy (anonymous):

and see what you will get

OpenStudy (anonymous):

Thank you!